OpenStudy (sammyalabamy):

I need some help with this problem. I know I need to use log, but I don't know how to set it up. Find the exact solution to the equation. Show your work for credit.

OpenStudy (sammyalabamy):

\[4^{(12-4x)}=256\]

OpenStudy (anonymous):

you don't need a log for this. you can write both sides with the same base. \[4^{(12-4x)}=4^4\]

OpenStudy (sammyalabamy):

okay thanks